# Aligning LLMs for FL-free Program Repair

### v1

@@ Abstract
Large language models (LLMs) have achieved decent results on automated program repair (APR). However, the next token prediction training objective of decoder-only LLMs (e.g., GPT-4) is misaligned with the masked span prediction objective of current infilling-style methods, which impedes LLMs from fully leveraging pre-trained knowledge for program repair. In addition, while some LLMs are capable of locating and repairing bugs end-to-end when using the related artifacts (e.g., test cases) as input, existing methods regard them as separate tasks and ask LLMs to generate patches at fixed locations. This restriction hinders LLMs from exploring potential patches beyond the given locations.
In this paper, we investigate a new approach to adapt LLMs to program repair. Our core insight is that LLM’s APR capability can be greatly improved by simply aligning the output to their training objective and allowing them to refine the whole program without first performing fault localization. Based on this insight, we designed D4C, a straightforward prompting framework for APR. D4C can repair 180 bugs correctly in Defects4J, with each patch being sampled only 10 times. This surpasses the SOTA APR methods with perfect fault localization by 10% and reduces the patch sampling number by 90%. Our findings reveal that (1) objective alignment is crucial for fully exploiting LLM’s pre-trained capability, and (2) replacing the traditional localize-then-repair workflow with direct debugging is more effective for LLM-based APR methods. Thus, we believe this paper introduces a new mindset for harnessing LLMs in APR.
@@ I Introduction
However, current application of LLMs is not well-adapted for program repair due to two major reasons. First, the objective of current approaches is misaligned with LLM’s training objective. Existing LLM-based methods [14, 18, 19, 15] typically adopt the infilling-style approach, i.e., an approach that predicts the masked span of sequence tokens given the preceding (prefix) and subsequent (suffix) code. These methods adopt the infilling style because its objective aligns with the objective of the encoder-only or encoder-decoder models used in their implementation. For example, AlphaRepair [14], one of the earliest LLM-based infilling methods, uses CodeBERT [20], an encoder-only model that is trained to denoise masked tokens. Recently, decoder-only LLMs (e.g., GPT-4 [21]) have shown stronger code comprehension capability [22], and recent research attempted to directly employ them as the patch generator for infilling-style APR [23, 24, 25], which demonstrate SOTA performance. However, as these decoder-only LLMs are trained to predict the next tokens given the past token context [13], there exists a misalignment between their training objective and the objective of infilling-style APR. Specifically, while decoder-only LLMs are proficient at generating a complete program because of the sufficient code corpus in pre-training, we hypothesize that these LLMs are less proficient in patch generation on masked buggy code due to the sparsity of parallel <masked, denoised> code corpus in pre-training. As noted in recent studies about LLM, objective misalignment can yield significant sub-optimal performance [26, 27].
@@
Second, the current workflow limits LLM from fully exploiting its pre-trained capability. Existing approaches still adhere to the workflow of first using fault localization (FL) tools [28] to obtain a ranked list of potential fault locations, and then using APR methods to generate patches by modifying the buggy program at these locations [2, 3]. However, LLMs already have the capability to fix trivial syntax bugs in their generated code independently by referring to different artifacts (e.g., error report from compilers and virtual machines) [15, 29, 30], illustrating the capability of identifying bugs without FL tools. Considering LLM’s superior code comprehension ability, and several studies show that current FL tools may fail to provide precise fault locations [31, 32], asking LLMs to generate patches at provided locations might contrarily hinder them from exploring broader potential patch space beyond the provided ones. A recent study also revealed that LLMs fail to make good use of the given buggy lines and tend to over-rely on them [17], whereas another study suggested using a more flexible fault localization to enhance APR [33]. As observed in these studies, the traditional localize-then-repair workflow may lead to ineffective LLM-based APR.
@@
In this work, we explore a novel way to use LLM for program repair. We propose that (1) aligning the output from infilling discrete hunks to completing entire functions can better attain the training objective, and (2) allowing LLM to locate and repair bugs with artifacts in a human-like manner can further improve its APR performance. Based on these insights, we implemented D4C (i.e., Direct Debug Drives Decent Code), a straightforward end-to-end APR approach without complex prompting or extra training. We conducted experiments on a total of 1027 bugs associated with individual functions (i.e., single-function bugs) from Defects4J [34] and DebugBench [35]. D4C repaired 180 out of the 437 single-function bugs in Defects4J, outperforming the state-of-the-art (SOTA) APR methods with perfect FL [24] by 10%. Meanwhile, each patch in D4C was sampled only 10 times, accounting for 10% used by SOTA methods (which use 100–5000 samples). Our results show that by fully aligning the output format and augmenting the input prompt, D4C can achieve the best end-to-end APR performance without FL, additional fine-tuning, or multi-round dialogue. This paper is not aimed at proposing D4C as a new APR technique, but rather to introduce a new mindset or paradigm for better harnessing LLMs in APR in the future.
@@
Instead of relying on fault localization to obtain the list of locations to modify as in traditional APR workflow, we show that allowing LLMs to end-to-end locate and repair bugs by using diverse types of artifacts (e.g., failed tests, error messages, and code comments) in a human-like manner can further improve its APR performance.
@@ II-A LLM Architectures and Training Objectives
Background: A large language model (LLM) is a language model consisting of a neural network with many parameters (typically billions of weights or more) trained on large quantities of unlabelled corpus using self-supervised learning [36]. The LLMs usually adopt the Transformer [37] architecture or its sub-structures (i.e., encoder or decoder). Based on their architectures, LLMs can be categorized into three types: encoder-only, decoder-only, and encoder-decoder LLMs. LLMs in the same architecture also share the same or similar training objectives. The details of the architecture and corresponding training objectives are shown in the Fig. 1.
Encoder-only LLMs, such as BERT [38] and its variants like CodeBERT [20], have a bidirectional transformer encoder structure. They are typically trained on the masked language modeling objective (i.e., MLM), aiming to denoise and reconstruct the masked tokens via understanding the surrounding context. As shown in Eq. 1, the loss of MLM training objective can be explicitly represented as the log-sum of the conditional probabilities of generating the masked token when all the unmasked tokens are known.
@@
Decoder-only LLMs, including GPT series [13, 21] and LLaMA series [39], have an autoregressive transformer decoder structure. They are mainly trained on the causal language modeling objective (i.e., CLM), aiming to predict and complete the next tokens via following the prefix context. As shown in Eq. 2, the loss of CLM training objective can be explicitly represented as the log-sum of the conditional probabilities of generating the next token when all the preceeding tokens are known.
@@
Encoder-decoder LLMs, such as T5 [40] and its variants like CodeT5 [41], have a complete transformer structure. They use an encoder to embed the input text into vector representations, and a decoder to causally generate new tokens after the input prompt. Specifically, T5 series LLMs are often pre-trained on the denoising objective, which aims to generate reconstructed spans (i.e., a sequence of adjacent tokens) for the masked span in the prompt. As shown in Eq. 3, the loss of denoising training objective can be explicitly represented as the log-sum of the conditional probabilities of generating the masked token when all input tokens are known.
@@
Motivation: As introduced, both the encoder-only and encoder-decoder (especially T5 series) LLMs involve the corruption and denoising process during pre-training, which substitutes some tokens in the input text for mask tokens, and reconstruct these tokens from the mask tokens. To make effective use of them for APR, Xia et al. [14]. proposed the infilling-style APR. It aims to replace the entire input buggy hunk with a mask token (e.g., <INFILL>) and prompts the model for direct inference, restoring the fixed hunk from the mask. The infilling-style APR can be implemented differently across various models. For instance, in CodeBERT, it recovers the correct code from masked hunks while leaving the rest of the input tokens the same [20]. Meanwhile, in CodeT5, it only outputs the correct code restored from the masked hunks, excluding all other input [41]. These task modeling strategies align with their model training objectives.
Figure 2: An example of current automated program repair workflow.
Figure 3: An example of two different APR paradigms. First row: Locate and repair subsequently may cause many invalid attempts when the localization is wrong. Second row: Locate and repair simultaneously can mitigate the limitation of patching at specific locations. (The wavy line denotes the fault location)
Recently, several APR approaches use decoder-only LLMs as they have exhibited a strong capacity for program comprehension and generation. Influenced by the prior use of infilling-style approaches, it has been generally assumed that such methods would be useful for APR. Thus, prior techniques use decoder-only LLMs to generate discrete fixed hunks for the given buggy hunks, as illustrated in Fig. 2. However, except for very few models (e.g., InCoder [42]), decoder-only LLMs barely incorporate denoising objective during pre-training. As the <masked, denoised> parallel corpus is rarely available, it becomes challenging for the models to learn the generation of the fixed hunk from the masked tokens without further adaptation after the next token prediction pre-training. Such an objective misalignment might hamper LLM’s performance on specific tasks extensively, such as the text classification task [27] and the QA task [26], whose task objectives are misaligned with the training objective of decoder-only LLMs. Thus, we have the following insight.
Insight 1: Modifying decoder-only LLMs’ output from discrete fixed hunks to the entire refined program can better align the inference objective to the training objective, thus significantly enhancing APR performance.
@@ II-B APR Techniques and Workflow
Background: Given a buggy program and an artifact (e.g., usually a test suite with at least one failed test), automated program repair (APR) approaches generate a fixed program that fulfills a correctness criteria (e.g., passing all tests). As illustrated in Fig. 2, a traditional APR workflow typically encompasses three steps: (1) fault localization, (2) patch generation, and (3) patch validation. Particularly, APR research mainly focuses on the patch generation step and obtains the fault locations using existing FL techniques (e.g., statistical fault localization) or uses perfect FL results in evaluation.
@@
Based on their patch generation strategies, APR methods can be categorized into heuristic-based [4, 5, 43, 44, 45, 46, 47], constraint-based [7, 48, 49], template-based [50, 8, 9], and learning-based  [51, 10, 11, 52, 53, 54, 55]. The non-learning-based approaches are usually restricted by a limited set of program transformations, causing the generation of a large number of invalid patches. To identify location to apply these transformations, the current strategy is to first generate a ranked list of suspicious fault locations using FL tools. The APR tool then sequentially generates patches for each provided fault location. Each of the generated patches is then validated using the given artifacts (e.g., test cases). With the advent of deep learning (DL), learning-based APR tools based on neural machine translation (NMT) have emerged. These tools outperform the traditional methods [3] via learning code semantics on a large bug-fix parallel training corpus without specially designed patch templates nor search heuristics. Moreover, since NMT models generally have a limited parameter size, their training costs and inference efficiency are manageable. Therefore, they are well-suited for the traditional APR workflow.
@@
Figure 4: The workflow of D4C. It uses the buggy code and its corresponding documents, failed tests, and test info (e.g., error message) to construct the prompt for one-shot prompting-based program repair without a specific fault location (usually provided by FL tools).
@@
Motivation: In recent years, there is a growing interest in using LLMs (e.g., GPT-4 [21]) for APR [14, 56, 33, 54]. Researchers usually follow the current APR workflow, positioning LLM as a new patch generator to replace former NMT models. However, taking LLM as a simple substitute in the patch generation step in the current workflow is an under-utilization of its pre-trained knowledge, since these models exhibit near-human capabilities in complex problem analysis and reasoning. For example, SELF-DEBUGGING has been proposed with the concept of allowing LLMs to progressively refine their generated code by engaging in multiple self-dialogue rounds using question explanations and execution traces, eventually producing high-quality code [30]. As shown in Fig. 3, since FL tools may not always provide perfect predictions of fault locations, restricting LLM to fixing the provided locations may lead to time waste in validating patches at incorrect locations. Moreover, as LLM’s inference overhead and time cost is much higher than traditional APR tools, using LLM to generate patches for all given locations may result in huge resource waste. Thus, we have the insight below:
@@
Insight 2: Prompting LLMs with buggy programs and corresponding artifacts (e.g., failing tests) can enable them to locate and repair bugs simultaneously without FL, thus further improving their APR performance.
@@
This insight arises from the APR methodology of programmers who typically identify bugs through tests and documents, and subsequently fix bugs based on these test results. Moreover, existing research revealed that models can produce decent fixes on basic compile bugs when supplemented with compiler errors [15], and their patch generation capability can be further improved with the assistant of the given failed test information [24, 30]. Thus, we believe this approach can overcome the limitation of generating patches within a limited number of locations, thereby further enhancing LLM’s APR performance. We will verify our hypothesis in Sec. IV-C via comparing the number of correct patches of two input formats (i.e., w/ or w/o artifacts) on Mixtral-MoE and GPT-4.

### v2

# Aligning the Objective of LLM-based Program Repair

@@ Abstract
Large language models (LLMs) have achieved decent results on automated program repair (APR). However, the next token prediction training objective of decoder-only LLMs (e.g., GPT-4) is misaligned with the masked span prediction objective of current infilling-style methods, which impedes LLMs from fully leveraging pre-trained knowledge for program repair. In addition, while some LLMs can locate and repair bugs in certain functions using the related artifacts (e.g., test cases), existing methods still depend on statement-level fault localization methods to provide a list of buggy hunks for repair. This restriction hinders LLMs from exploring potential patches beyond the given locations.
In this paper, we investigate a new approach to adapt LLMs to program repair. Our core insight is that LLM’s APR capability can be greatly improved by simply aligning the output to their training objective and allowing them to refine the whole program without first identifying faulty statements. Based on this insight, we designed D4C, a straightforward prompting framework for APR. D4C can repair 180 bugs correctly in Defects4J, with each patch being sampled only 10 times. This surpasses the SOTA APR methods with perfect fault localization by 10% and reduces the patch sampling number by 90%. Our findings reveal that (1) objective alignment is crucial for fully exploiting LLM’s pre-trained capability, and (2) replacing the traditional localize-buggy-hunks-then-repair workflow with direct debugging is more effective for LLM-based APR methods. Thus, we believe this paper introduces a new mindset for harnessing LLMs in APR.
Index Terms:
Automated Program Repair, Large Language Model, Objective Alignment
@@ I Introduction
However, current application of LLMs is not well-adapted for program repair due to two major reasons. First, the inference objective of current approaches is misaligned with LLM’s training objective. LLMs can be classified into encoder-only, encoder-decoder, and decoder-only. The first two are typically trained to predict the masked tokens (i.e., infilling [18] or denoising [19]), while the third is trained to predict the next tokens (i.e., completion [13]). Existing LLM-based methods [14, 20, 21] generally adopt the infilling-style method [14] to predict the fixed code at the masked buggy code location during inference. The early work achieved decent results, as the models used are also trained on infilling task [14, 20]. Recently, larger decoder-only LLMs like GPT-4 [22] have shown stronger capability on code tasks [23], and people attempt to directly employ them for infilling-style APR [24, 25, 26]. However, while the model parameters being scaled up by hundreds of times, the fixed bugs did not even doubled [14, 25], which contrasts with the clear scalability in other tasks [23]. We hypothesize that this is due to decoder-only LLMs are trained for completion rather than infilling, resulting in an objective misalignment between training and inference. Specifically, due to the sparsity of parallel <masked, denoised> code corpus in training, these LLMs are less proficient in patch generation on masked buggy code. As noted in recent studies [27, 28], objective misalignment can yield significant sub-optimal performance.
@@
Second, the current workflow limits LLM from fully exploiting its pre-trained capability. Existing approaches still adhere to the workflow of first using statement-level fault localization (statement-level FL) tools [29] to obtain a ranked list of potential buggy hunks (i.e., contiguous statements to be modified in patches [30]), and then using APR methods to generate patches by modifying the buggy program at these hunks [2, 3]. However, LLMs already have the capability to fix trivial syntax bugs in their generated code independently by referring to different artifacts (e.g., error report from compilers and virtual machines) [15, 31, 32], illustrating the capability of identifying bugs without given buggy hunks. Considering LLM’s superior code comprehension ability, and several studies show that current FL tools may fail to provide precise buggy hunks [33, 34], asking LLMs to generate patches at provided several hunks might contrarily hinder them from exploring broader potential patch space beyond the provided ones. A recent study also revealed that LLMs fail to make good use of the given buggy lines and tend to over-rely on them [17], whereas another study suggested using a more flexible fault localization to enhance APR [35]. As observed in these studies, the traditional localize-buggy-hunks-then-repair workflow may lead to ineffective LLM-based APR.
@@
In this work, we explore a novel way to use LLM for program repair. We propose that (1) aligning the output from infilling discrete hunks to completing entire functions can better attain the training objective, and (2) allowing LLM to locate and repair buggy hunks with artifacts in a human-like manner can further improve its APR performance. Based on these insights, we implemented D4C (i.e., Direct Debug Drives Decent Code), a straightforward APR approach without complex prompting or extra training. We conducted experiments on a total of 1027 bugs associated with individual functions (i.e., single-function bugs) from Defects4J [36] and DebugBench [37]. D4C repaired 180 out of the 437 single-function bugs in Defects4J, outperforming the state-of-the-art (SOTA) APR methods with perfect statement-level FL [25] by 10%. Meanwhile, each patch in D4C was sampled only 10 times, accounting for 10% used by SOTA methods (which use 100–5000 samples). Our results show that by fully aligning the output format and augmenting the input prompt, D4C can achieve the best APR performance without given buggy hunks, additional fine-tuning, or multi-round dialogue. This paper is not aimed at proposing D4C as a new APR technique, but rather to introduce a new mindset or paradigm for better harnessing LLMs in APR in the future.
@@
Instead of relying on statement-level FL tools to obtain the list of buggy hunks for repair as in traditional APR workflow, we show that allowing LLMs to locate and repair buggy hunks by using diverse types of artifacts (e.g., failed tests, error messages, and code comments) in a human-like manner can further improve its APR performance.
@@ II-A LLM Architectures and Training Objectives
Background: A large language model (LLM) is a language model consisting of a neural network with many parameters (typically billions of weights or more) trained on large quantities of unlabelled corpus using self-supervised learning [38]. The LLMs usually adopt the Transformer [39] architecture or one of its sub-structures (i.e., encoder or decoder). The encoder usually consists of feed-forward networks with self-attention [39], while the decoder usually consists of feed-forward networks with cross-attention [39]. Thus, LLMs can be categorized into three types: encoder-only, decoder-only, and encoder-decoder LLMs.
Encoder-only LLMs, such as BERT [18] and its variants like CodeBERT [40], have a bidirectional transformer encoder structure. They are typically trained on the masked language modeling objective (i.e., MLM), aiming to denoise and reconstruct the masked tokens via understanding the surrounding context (Fig. 1). As shown in Eq. 1, the loss of MLM training objective can be explicitly represented as the log-sum of the conditional probabilities of generating the masked token when all the unmasked tokens are known.
@@
Decoder-only LLMs, including GPT series [13, 22] and LLaMA series [41], have an autoregressive transformer decoder structure. They are mainly trained on the causal language modeling objective (i.e., CLM), aiming to predict and complete next tokens via following the prefix context (Fig. 1). As shown in Eq. 2, the loss of CLM training objective can be explicitly represented as the log-sum of the conditional probabilities of generating the next token when all the preceeding tokens are known.
@@
Encoder-decoder LLMs, such as T5 [19] and its variants like CodeT5 [42], have a complete transformer structure. They use an encoder to embed the input text into vector representations, and a decoder to causally generate new tokens after the input prompt. Specifically, T5 series LLMs are often pre-trained on the denoising objective, which aims to generate the reconstructed spans (i.e., a sequence of adjacent tokens) from the masked spans for the masked span in the prompt (Fig. 1). As shown in Eq. 3, the loss of denoising training objective can be explicitly represented as the log-sum of the conditional probabilities of generating the masked token when all input tokens are known.
@@
Figure 2: An example of automated program repair workflow.
Motivation: As introduced, both the encoder-only and encoder-decoder (especially T5 series) LLMs involve the corruption and denoising process during pre-training, which substitutes some tokens in the input text for mask tokens, and reconstruct these tokens from the mask tokens. To make effective use of them for APR, Xia et al. [14] proposed the infilling-style APR. It aims to replace the entire input buggy hunk with a mask token (e.g., <INFILL>) and prompts the model for direct inference, restoring the fixed hunk from the mask. The infilling-style APR can be implemented differently across various models. For instance, in CodeBERT, it recovers the correct code from masked hunks while leaving the rest of the input tokens the same [40]. In CodeT5, it only outputs the correct code restored from the masked hunks [42].
Figure 3: An example of two different APR paradigms. First row: Locate and repair buggy hunks subsequently may cause many invalid attempts for patch generation. Second row: Locate and repair buggy hunks simultaneously can mitigate the cost of patching at specific hunks. (The wavy line is a buggy hunk)
Recently, several APR approaches use decoder-only LLMs as they have exhibited a stronger coding capacity. Influenced by the prior success of infilling, it has been generally assumed that such methods would be useful for decoder-only LLMs. Thus, they use decoder-only LLMs to generate fixed hunks for the given buggy hunks, as illustrated in Fig. 2. However, except for very few models (e.g., InCoder [43]), decoder-only LLMs barely incorporate denoising objective during pre-training. As the <masked, denoised> parallel corpus is rarely available, it becomes challenging for the models to learn the generation of the fixed hunk from the masked tokens without further adaptation after the next token prediction pre-training. Such an objective misalignment might hamper LLM’s performance on specific tasks extensively, such as the text classification task [28] and the QA task [27], whose inference objectives are misaligned with the training objective of decoder-only LLMs. Thus, we have the following insight.
Insight 1: Modifying decoder-only LLMs’ output from fixed hunks to the entire refined program can better align the inference objective to the training objective, thus significantly enhancing APR performance.
@@ II-B APR Techniques and Workflow
Background: Given a buggy program and an artifact (e.g., usually a test suite with at least one failed test), automated program repair (APR) approaches generate a fixed program that fulfills a correctness criteria (e.g., passing all tests). As illustrated in Fig. 2, an APR workflow typically encompasses three steps: (1) fault localization, (2) patch generation, and (3) patch validation. Particularly, APR research mainly focuses on the patch generation step and obtains the fault locations using existing FL techniques (e.g., statistical fault localization) or uses perfect FL results in evaluation.
@@
Based on their patch generation strategies, APR methods can be categorized into heuristic-based [4, 5, 44, 45, 46, 47, 48], constraint-based [7, 49, 50], template-based [51, 8, 9], and learning-based [52, 10, 11, 53, 54, 55, 56]. The non-learning-based approaches are usually restricted by a limited set of program transformations, causing the generation of a large number of invalid patches. To identify location to apply these transformations, the current strategy is to first generate a ranked list of suspicious buggy hunks using statement-level FL tools. The APR tool then sequentially generates patches for each provided hunks. Each of the generated patches is then validated using the given artifacts (e.g., test cases). With the advent of deep learning (DL), learning-based APR tools based on neural machine translation (NMT) have emerged. These tools outperform the traditional methods [3] via learning code semantics on a large bug-fix parallel training corpus without specially designed patch templates nor search heuristics. Moreover, since NMT models generally have a limited parameter size, their training costs and inference efficiency are manageable. Therefore, they are well-suited for the traditional APR workflow.
@@
Figure 4: The workflow of D4C. It uses the buggy code and its corresponding documents, failed tests, and test info (e.g., error message) to construct the prompt for one-shot prompting-based program repair without a specific buggy hunk (usually provided by statement-level FL tools).
@@
Motivation: In recent years, there is a growing interest in using LLMs (e.g., GPT-4 [22]) for APR [14, 57, 35, 55]. Researchers usually follow the current APR workflow, positioning LLM as a new patch generator to replace former NMT models. However, taking LLM as a simple substitute in the patch generation step in the current workflow is an under-utilization of its pre-trained knowledge, since these models exhibit the capability of locating and fixing bugs in buggy functions independently. For example, SELF-DEBUGGING [32] has been proposed with the concept of allowing LLMs to progressively refine their generated buggy functions to produce bug-free functions by engaging in multiple self-dialogue rounds with artifacts like execution traces, without using statement-level FL to identify buggy hunks. As shown in Fig. 3, since statement-level FL tools may not always provide perfect predictions of buggy hunks, restricting LLM to fixing the provided hunks may lead to time waste in validating patches at incorrect locations. Moreover, as LLM’s inference overhead and time cost is much higher than traditional APR tools, using LLM to generate patches for all given hunks may result in huge resource waste. Thus, we have the insight below:
@@
Insight 2: Prompting LLMs with buggy programs and corresponding artifacts can enable them to locate and repair buggy hunks simultaneously without statement-level FL, thus further improving APR performance.
@@
This insight arises from the APR methodology of programmers who typically identify buggy hunks through tests and documents, and subsequently fix bugs based on these test results. Moreover, existing research revealed that models can produce decent fixes on basic compile bugs when supplemented with compiler errors [15], and their patch generation capability can be further improved with the assistant of the given failed test information [25, 32]. Thus, we can use a more flexible FL method [35] to identify the buggy segments of the program at a coarse-grained level (e.g., using method-level FL [58] to find the buggy function), and allow the LLM to simultaneously locate and fix the buggy hunks within those segments, as locating the segments containing the buggy hunks is easier and more practical than using statement-level FL to directly find the correct buggy hunks for repair. We will verify our hypothesis in Sec. IV-C via comparing the number of correct patches of two input formats (i.e., w/ or w/o artifacts).